2. ABC Corp. issued 15-year bonds with a par value of $1,000 2 years ago at a coupon rate of 5.9 percent. The bonds make semiannual payments. If these bonds currently sell for $1,050, what is the YTM?
Answer #1
YTM = =rate(nper,pmt,-pv,fv)*2
= 5.36%
Where,
nper = (15-2)*2 = 26
pmt = 1000*5.9%*1/2 = $       29.50
pv = $ 1,050.00
fv = $ 1,000.00